a and b are positive integers. Find 2a+3b if a^2b^3=108
Answer by fractalier(6550)

Given a^2b^3=108, the only way this can be is if a = 2 and b = 3...because
2^2 * 3^3 = 4 * 27 = 108...if so, then
2a + 3b = 2*2 + 3*3 = 13